# Parton Distribution Function of a Deuteron-like Dibaryon System from Lattice QCD

### v1

@@ Abstract
We report a lattice QCD calculation of the parton distribution function (PDF) of a deuteron-like dibaryon system using large-momentum effective theory. The calculation is done on three Wilson Clover ensembles with a fixed lattice spacing $a=0.105$ fm and two pion masses. The lattice matrix elements are computed at proton momenta up to 2.46 GeV with the signal of high momentum modes being improved by applying the momentum smearing technique. The state-of-the-art renormalization, matching and extrapolation are then applied to obtain the final result of the light-cone PDF. A comparison between the result of the dibaryon system and the sum of the proton and neutron PDFs is also given.
@@ I Introduction
Parton distribution functions (PDFs) are widely applied in high-energy physics. On the one hand, they are necessary inputs for computing cross sections in electron-hadron or hadron-hadron collisions. On the other hand, they contain important information about the intrinsic properties of nucleons and nuclei. To date, the PDFs of single-baryon systems like the proton or neutron have been extensively studied both experimentally and theoretically. In the meantime, the PDFs of multi-baryon systems like the deuteron and 3He have also attracted a lot of attention. Comparing the structure functions between single-baryon and multi-baryon systems reveals important insights into the interaction between baryons within a nucleus. For example, the European Muon Collaboration (EMC) first observed in 1983 that the structure functions of the nucleus differ from the sum of the structure functions of individual nucleons inside the nucleus Aubert et al. (1983), which is now known as the“nuclear EMC effect”. It can be characterized by the ratio
@@
with $F_{2}(x)=x\sum\limits_{i}f_{i}(x)Q_{i}^{2}$ at the leading-order, $f_{i}(x)$ is the PDF of the $i$th quark flavor. The observation of the EMC effect has opened up new avenues of research in nuclear physics, leading to extensive further study and investigation through many experiments Ashman et al. (1988); Bodek et al. (1983); Accardi et al. (2016); Cocuzza et al. (2021); Segarra et al. (2021). On the theoretical side, substantial efforts have also been made to constrain nucleon-nucleon interactions by calculating basic properties of multi-nucleon systems using the lattice QCD (LQCD) approach Amarasinghe et al. (2023); Detmold et al. (2021); Illa (2022); Beane et al. (2006, 2007, 2011a, 2011b, 2012, 2013); Orginos et al. (2015); Wagman et al. (2017); Fukugita et al. (1995); Yamazaki et al. (2011, 2012, 2015); Nemura et al. (2009); Inoue (2012); Berkowitz et al. (2017); Francis et al. (2019); Junnarkar and Mathur (2019); Hörz et al. (2021); Illa et al. (2021); Green et al. (2021); Iritani et al. (2019); Murakami et al. (2023). Among them, the NPLQCD group has pioneered the calculation of moments of the proton, di-proton and 3He systems Detmold et al. (2021); Illa (2022).
@@
In addition, recent theoretical developments Braun and Müller (2008); Ji (2013, 2014); Ma and Qiu (2018); Lin et al. (2018a); Radyushkin (2017) have also enabled the direct calculation of the Bjorken $x$-dependence of PDFs from LQCD. In the past few years, various LQCD calculations for nucleon PDFs have been done Liu et al. (2020); Chen et al. (2018); Lin et al. (2018b, c); Liu et al. (2018); Yao et al. (2023a); Orginos et al. (2017); Alexandrou et al. (2015, 2020); Gao et al. (2023); Karpie et al. (2021); Edwards et al. (2023), based on either the large-momentum effective theory(LaMET) Ji (2013, 2014); Ji et al. (2021a) or the short-distance expansion (pseudo-PDF) approach Radyushkin (2017). Applying such calculations to multi-baryon systems will help further elucidate the origin of nuclear effects.
@@
The deuteron is the simplest multi-baryon bound system. It is composed of a proton and a neutron and has a small binding energy of $2.224644(34)$ MeV. Such a low binding energy poses a significant challenge for identifying a bound state deuteron through LQCD calculations of two-point correlation functions. However, utilizing three-point correlation functions, such as those employed in the calculation of PDFs, could potentially serve as a useful tool to study nuclear effects,as revealed by the EMC effect.
@@
In this work, we calculate the PDF of a deuteron-like dibaryon system using the LaMET approach. The ultraviolet (UV) divergences are removed by a nonperturbative renormalization in the hybrid scheme Ji et al. (2021b), which has been widely used in previous studies Yao et al. (2023a); Hua et al. (2022, 2021). While only one single lattice spacing is considered in this work, we perform our calculations on ensembles with two different pion masses at $941$ MeV and $293$ MeV. To assess the impact of finite volume effects, we conduct simulations at the lower pion mass of $293$ MeV on two sets of ensembles with lattice size $32^{3}\times 64$ and $24^{3}\times 72$. In order to study the nuclear effect, we make a comparison between the PDF of the dibaryon system and the sum of the proton and neutron PDFs.
@@ II.1 Lattice setup
Throughout the paper, we use the $2+1$ flavor QCD ensembles with stout smeared clover fermion action and Symanzik gauge actions generated by the CLQCD collaboration Hu et al. (2023); Du et al. (2024). Three enembles, named C32P29, C24P29 and C24P90 are used to calculate the bare matrix elements of the nucleon and dibaryon systems. The detailed information is given in Table 1.
@@
II.2 Light-cone PDF v.s. Quasi-PDF
@@
where $|P\rangle$ denotes the ground state of a nucleon or a multi-nucleon system with momentum $P$, $x$ is the momentum fraction carried by the quark, $\mu$ is the renormalization scale and $\xi^{\pm}=(\xi^{t}\pm\xi^{z})/\sqrt{2}$ are light-cone coordinates. The Wilson line along the light-cone direction
@@
is introduced to ensure gauge invariance of the non-local quark bilinear correlator.
The PDF involves real-time dependence and cannot be directly accessed on the lattice. However, according to LaMET, one can start from the calculation the quasi-PDF, which takes the following form
@@
where the matrix element on the r.h.s. is often called the quasi-light-front (quasi-LF) correlation, $P_{z}$ representing the momentum of a nucleon or a multi-nucleon system along $z$-direction. $\Gamma$ can be chosen as $\gamma_{t}$ or $\gamma_{z}$ for unpolarized quark quasi-PDFs. In this work, we choose $\Gamma=\gamma_{t}$, which is free from operator mixing with scalar quark quasi-PDF operator Constantinou and Panagopoulos (2017); Green et al. (2018); Chen et al. (2019). The relation between the quasi-PDF and the light-cone PDF is given by a factorization formula
@@
where $q(y,\mu)$ is the light-cone PDF depending on the renormalization scale, $C\left(\frac{x}{y},\frac{\mu}{yP_{z}}\right)$ is the matching kernel and $\mathcal{O}\left(\frac{\Lambda_{\mathrm{QCD}}^{2}}{\left(xP_{z}\right)^{2}},\frac{\Lambda_{\mathrm{QCD}}^{2}}{\left((1-x)P_{z}\right)^{2}}\right)$ denote higher-twist contributions suppressed by $P_{z}$.
@@ II.3 Two-point and three-point correlators
Here $N_{\sigma}(\vec{P},t)$ denotes the nucleon or multi-nucleon interpolating operators projected to certain momentum $\vec{P}$; $t_{sep}$ denotes the time interval between the source and the sink with the source being placed at $t=0$, and $\sigma$ labels the rows of a certain irreducible representation under which the operator transforms. $O_{\Gamma}(z;t)=\bar{\psi}(z;t)\Gamma U(z,0;t)\psi(0;t)$ where $\Gamma=\gamma_{t}$ is the current operator located at time $t$ with quark and anti-quark separated along the $z$ direction.
We construct the interpolating operators of nucleon and deuteron following the method proposed in Ref. Amarasinghe et al. (2023). Since a nucleon has quantum numbers $J^{P}=\frac{1}{2}^{+}$, the operator should transform in the $G_{1}^{+}$ irrep of the cubic group. The proton interpolating operator can be written as
@@
where quark fields are labeled with indices $i,j,k,...,$ which are a combination of spinor and color indices. For example, $i=(\zeta,a)$ with spinor index $\zeta\in\{0,1,2,3\}$ and color index $a\in\{0,1,2\}$; $w_{\alpha}^{[N]\sigma}$ denotes the weights of different spin-color combination, with $\alpha\in\{1,...,{\cal{N}}_{\omega}^{\left[N\right]}\}$ running over the ${\cal{N}}_{\omega}^{\left[N\right]}$ combinations of $u^{i\left(\alpha\right)}d^{j\left(\alpha\right)}u^{k\left(\alpha\right)}$. For each $\sigma$, only $12$ terms appear due to the sparseness of gamma matrices and antisymmetric tensor $\epsilon^{abc}$, which significantly reduces the computation time for contractions.
@@
In our calculation $\vec{P}=(0,0,(2\pi n)/{L})$ is directed along the $z$-axis. In principle, the operators should be summed over all spatial lattice sites for both the sink and the source. However, this approach would necessitate the calculation of all-to-all propagators, which is infeasible given our limited computational resources. As a practical compromise, we sum over all spatial lattice sites for the sink and a few points evenly distributed in the $z=0$ plane for the source. Specifically, we contract point-to-all propagators with sources fixed at one point, enhance the measurements by shifting the source position, and repeat the calculation. This method is also applied to the calculation of deuteron-like dibaryon correlators.
@@
where the subscripts of $D_{\{1,2,3\}}$ represent the three components of $J=1$, and $\vec{P}$ is the center of mass momentum of the system. For simplicity we only consider the case in which both the proton and the neutron carry half of the total momentum. $\vec{x}_{1}$ and $\vec{x}_{2}$ are space coordinates of the two nucleons. These operators are used at the sink, while at the source the two nucleons are put at the same space position since we compute point-to-all quark propagators as explained above.
The quark-level representation of dibaryon operators can be constructed by inserting Eq. (8) into Eq. (10):
@@
where the weights $w_{\alpha}^{[D]\rho}$, with $\alpha\in\{1,\ldots,\mathcal{N}_{w}^{[D]\rho}\}$, with $\mathcal{N}_{w}^{[D]1}=\mathcal{N}_{w}^{[D]3}=144$ and $\mathcal{N}_{w}^{[D]2}=288$, are obtained from products of weights of nucleon $w_{\alpha}^{[N]\sigma}$ as shown in  Amarasinghe et al. (2023).
The three-point correlation function can be obtained by summing over all possible contractions of connected diagrams. Our calculation did not account for the disconnected diagram contributions, which are not expected to have a significant impact our qualitative findings Alexandrou et al. (2021). In Fig. 1 we give an illustration of one possible contraction of the three-point correlator of the deuteron-like dibaryon system.
Figure 1: Illustration of the three-point correlator of a deuteron-like system before momentum projection (one possible way of contraction).
Three ensembles, namely C32P29, C24P29 and C24P90, are used to calculate two-point and three-point correlators of the deuteron-like dibaryon system and the nucleon. The calculations are performed with the nucleon carrying different momenta $P_{z}^{nucleon}=(2\pi n)/L$ along the $z$-direction and the total momentum of the dibaryon system is $2P_{z}^{nucleon}$ For each momentum, several sets of source-sink separations $t_{sep}$ are calculated. For larger momenta, correlators are averaged over several sources evenly distributed on the $z=0$ plane to control the statistical uncertainties. The detailed information is shown in Table 2.
Ensemble $P_{z}^{nucleon}$(GeV) $t_{sep}/a$ $N_{conf}$ $nsrc(time\times space)$
@@
Table 2: Information on neucleon’s momentum $P_{z}^{nucleon}$ (the center-of-mass momentum of dibaryon system is $2P_{z}^{nucleon}$), time separation between source and sink($t_{sep}$) and number of sources($nsrc$) on the $z=0$ plane used to calculate two-point and three-point correlators.
To improve the signal of high momentum states, we applied momentum-Gaussian smearing on sources and sinks Bali et al. (2016), which takes the form
@@
where $k$ represents the momentum-smearing parameter, $U_{j}$ denotes a stout-smeared gauge link along a specified direction. The above operation should be iterated by $n_{iter}=\mathcal{O}(2\sigma^{2})$ times to achieve the desired smearing size $\sigma a$, which corresponds to $n_{iter}=50$ and $\sigma a=0.5$ fm in our calculation. This process shifts the smearing center to momentum $\mathcal{O}(k)$, thereby significantly improving the signal of high-momentum states. In our calculation we set $k$ equals to half of the nucleon momentum , which is the optimized choice to improve the signal according to previous experience. In addition, we applied one-step hypercubic(HYP)-smearing on the Wilson link of the current operator.

### v2

# Parton distribution function of a deuteronlike dibaryon system from lattice QCD

@@ Abstract
We report a lattice QCD calculation of the parton distribution function (PDF) of a deuteronlike dibaryon system using large-momentum effective theory. The calculation is done on three Wilson-clover ensembles with a fixed lattice spacing $a=0.105$ fm and two pion masses. The lattice matrix elements are computed at proton momenta up to 2.46 GeV, with the signal of high-momentum modes being improved by applying the momentum smearing technique. The state-of-the-art renormalization, matching, and extrapolation are then applied to obtain the final result of the light-cone PDF. A comparison between the result of the dibaryon system and the sum of the proton and neutron PDFs is also given.
@@ I Introduction
Parton distribution functions (PDFs) are widely applied in high-energy physics. On the one hand, they are necessary inputs for computing cross sections in electron-hadron or hadron-hadron collisions. On the other hand, they contain important information about the intrinsic properties of nucleons and nuclei. To date, the PDFs of single-baryon systems like the proton or neutron have been extensively studied both experimentally and theoretically. In the meantime, the PDFs of multibaryon systems like the deuteron and 3He have also attracted a lot of attention. Comparing the structure functions between single-baryon and multibaryon systems reveals important insights into the interaction between baryons within a nucleus. For example, the European Muon Collaboration (EMC) first observed in 1983 that the structure functions of the nucleus differ from the sum of the structure functions of individual nucleons inside the nucleus Aubert et al. (1983), which is now known as the “nuclear EMC effect.” It can be characterized by the ratio
@@
with $F_{2}(x)=x\sum\limits_{i}f_{i}(x)Q_{i}^{2}$ at the leading order, where $f_{i}(x)$ is the PDF of the $i$-th quark flavor. The observation of the EMC effect has opened up new avenues of research in nuclear physics, leading to extensive further study and investigation through many experiments Ashman et al. (1988); Bodek et al. (1983); Accardi et al. (2016); Cocuzza et al. (2021); Segarra et al. (2021). On the theoretical side, substantial efforts have also been made to constrain nucleon-nucleon interactions by calculating basic properties of multinucleon systems using the lattice QCD (LQCD) approach Amarasinghe et al. (2023); Detmold et al. (2021); Illa (2022); Beane et al. (2006, 2007, 2011a, 2011b, 2012, 2013); Orginos et al. (2015); Wagman et al. (2017); Fukugita et al. (1995); Yamazaki et al. (2011, 2012, 2015); Nemura et al. (2009); Inoue (2012); Berkowitz et al. (2017); Francis et al. (2019); Junnarkar and Mathur (2019); Hörz et al. (2021); Illa et al. (2021); Green et al. (2021); Iritani et al. (2019); Murakami et al. (2023). Among them, the NPLQCD group has pioneered the calculation of moments of the proton, diproton, and 3He systems Detmold et al. (2021); Illa (2022).
@@
In addition, recent theoretical developments Braun and Müller (2008); Ji (2013, 2014); Ma and Qiu (2018); Lin et al. (2018a); Radyushkin (2017) have also enabled the direct calculation of the Bjorken $x$-dependence of PDFs from LQCD. In the past few years, various LQCD calculations for nucleon PDFs have been done Liu et al. (2020); Chen et al. (2018); Lin et al. (2018b, c); Liu et al. (2018); Yao et al. (2023a); Orginos et al. (2017); Alexandrou et al. (2015, 2020); Gao et al. (2023); Karpie et al. (2021); Edwards et al. (2023), based on either the large-momentum effective theory(LaMET) Ji (2013, 2014); Ji et al. (2021a) or the short-distance expansion (pseudo-PDF) approach Radyushkin (2017). Applying such calculations to multibaryon systems will help to further elucidate the origin of nuclear effects.
@@
The deuteron is the simplest multibaryon bound system. It is composed of a proton and a neutron and has a small binding energy of $2.224575(9)$ MeV Van Der Leun and Alderliesten (1982). Such a low binding energy poses a significant challenge for identifying a bound-state deuteron through LQCD calculations of two-point correlation functions. Utilizing three-point correlation functions, such as those employed in the calculation of PDFs faces the same problem. However, it could still serve as a potentially useful and complementary tool to study nuclear effects, as revealed by the EMC effect.
@@
In this work, we calculate the PDF of a deuteronlike dibaryon system using the LaMET approach. The ultraviolet (UV) divergences are removed by a nonperturbative renormalization in the hybrid scheme Ji et al. (2021b), which has been widely used in previous studies Yao et al. (2023a); Hua et al. (2022, 2021). While only one single lattice spacing is considered in this work, we perform our calculations on ensembles with two different pion masses at $941$ MeV and $293$ MeV. To assess the impact of finite-volume effects, we conduct simulations at the lower pion mass of $293$ MeV on two sets of ensembles with lattice sizes of $32^{3}\times 64$ and $24^{3}\times 72$. In order to study the nuclear effect, we make a comparison between the PDF of the dibaryon system and the sum of the proton and neutron PDFs.
@@ II.1 Lattice setup
Throughout the paper, we use the $2+1$ flavor QCD ensembles with stout-smeared clover fermion action and Symanzik gauge actions generated by CLQCD collaboration Hu et al. (2023); Du et al. (2024). Three ensembles, named C32P29, C24P29 and C24P90 are used to calculate the bare matrix elements of the nucleon and dibaryon systems. The detailed information is given in Table 1.
@@
II.2 Light-cone PDF vs quasi-PDF
@@
where $|P\rangle$ denotes the ground state of a nucleon or a multinucleon system with momentum $P$, $x$ is the momentum fraction carried by the quark, $\mu$ is the renormalization scale and $\xi^{\pm}=(\xi^{t}\pm\xi^{z})/\sqrt{2}$ are light-cone coordinates. The Wilson line along the light-cone direction
@@
is introduced to ensure gauge invariance of the nonlocal quark bilinear correlator.
The PDF involves real-time dependence and cannot be directly accessed on the lattice. However, according to LaMET, one can start from the calculation of the quasi-PDF, which takes the following form:
@@
where the matrix element on the rhs is often called the quasi-light-front (quasi-LF) correlation. $P_{z}$ represents the momentum of a nucleon or a multinucleon system along the $z$ direction. $\Gamma$ can be chosen as $\gamma_{t}$ or $\gamma_{z}$ for unpolarized quark quasi-PDFs. In this work, we choose $\Gamma=\gamma_{t}$, which is free from operator mixing with a scalar quark quasi-PDF operator Constantinou and Panagopoulos (2017); Green et al. (2018); Chen et al. (2019). The relation between the quasi-PDF and the light-cone PDF is given by a factorization formula
@@
where $q(y,\mu)$ is the light-cone PDF depending on the renormalization scale, $C\left(\frac{x}{y},\frac{\mu}{yP_{z}}\right)$ is the matching kernel relevant to the longitudinal partonic momentum  Izubuchi et al. (2018), and $\mathcal{O}\left(\frac{\Lambda_{\mathrm{QCD}}^{2}}{\left(xP_{z}\right)^{2}},\frac{\Lambda_{\mathrm{QCD}}^{2}}{\left((1-x)P_{z}\right)^{2}}\right)$ denote higher-twist contributions suppressed by $P_{z}$.
@@ II.3 Two-point and three-point correlators
Here, $N_{\sigma}(\vec{P},t)$ denotes the nucleon or multinucleon interpolating operators projected to certain momentum $\vec{P}$, $t_{sep}$ denotes the time interval between the source and the sink, with the source being placed at $t=0$, and $\sigma$ labels the rows of a certain irreducible representation, under which the operator transforms. $O_{\Gamma}(z;t)=\bar{\psi}(z;t)\Gamma U(z,0;t)\psi(0;t)$, where $\Gamma=\gamma_{t}$ is the current operator located at time $t$ with quark and antiquark separated along the $z$ direction.
We construct the interpolating operators of the nucleon and deuteron following the method proposed in Ref. Amarasinghe et al. (2023). Since a nucleon has quantum numbers $J^{P}=\frac{1}{2}^{+}$, the operator should transform in the $G_{1}^{+}$ irrep of the cubic group. The proton interpolating operator can be written as
@@
where quark fields are labeled with indices $i,j,k,...,$ which are a combination of spinor and color indices. For example, $i=(\zeta,a)$ with spinor index $\zeta\in\{0,1,2,3\}$ and color index $a\in\{0,1,2\}$, and $w_{\alpha}^{[N]\sigma}$ denotes the weights of different spin-color combinations, with $\alpha\in\{1,...,{\cal{N}}_{\omega}^{\left[N\right]}\}$ running over the ${\cal{N}}_{\omega}^{\left[N\right]}$ combinations of $u^{i\left(\alpha\right)}d^{j\left(\alpha\right)}u^{k\left(\alpha\right)}$. For each $\sigma$, only $12$ terms appear due to the sparseness of gamma matrices and antisymmetric tensor $\epsilon^{abc}$, which significantly reduces the computation time for contractions.
@@
In our calculation, $\vec{P}=(0,0,(2\pi n)/{L})$ is directed along the $z$-axis. In principle, the operators should be summed over all spatial lattice sites for both the sink and the source. However, this approach would necessitate the calculation of all-to-all propagators, which is infeasible given our limited computational resources. As a practical compromise, we sum over all spatial lattice sites for the sink and a few points evenly distributed in the $z=0$ plane for the source. Specifically, we contract point-to-all propagators with sources fixed at one point, enhance the measurements by shifting the source position, and repeat the calculation. This method is also applied to the calculation of deuteronlike dibaryon correlators.
@@
where the subscripts of $D_{\{1,2,3\}}$ represent the three components of $J=1$, and $\vec{P}$ is the center-of-mass momentum of the system. For simplicity, we only consider the case in which both the proton and the neutron carry half of the total momentum. $\vec{x}_{1}$ and $\vec{x}_{2}$ are space coordinates of the two nucleons. These operators are used at the sink, while at the source, the two nucleons are put at the same space position since we compute point-to-all quark propagators, as explained above.
The quark-level representation of dibaryon operators can be constructed by inserting Eq. (8) into Eq. (10),
@@
where the weights $w_{\alpha}^{[D]\rho}$, with $\alpha\in\{1,\ldots,\mathcal{N}_{w}^{[D]\rho}\}$, with $\mathcal{N}_{w}^{[D]1}=\mathcal{N}_{w}^{[D]3}=144$ and $\mathcal{N}_{w}^{[D]2}=288$, are obtained from products of weights of nucleon $w_{\alpha}^{[N]\sigma}$, as shown in  Amarasinghe et al. (2023).
The three-point correlation function can be obtained by summing over all possible contractions of connected diagrams. Our calculation does not account for the disconnected diagram contributions, which are not expected to have a significant impact on our qualitative findings Alexandrou et al. (2021). In Fig. 1 we give an illustration of one possible contraction of the three-point correlator of the deuteronlike dibaryon system.
Figure 1: Illustration of the three-point correlator of a deuteronlike system before momentum projection (one possible way of contraction).
Three ensembles, namely, C32P29, C24P29, and C24P90, are used to calculate two-point and three-point correlators of the deuteronlike dibaryon system and the nucleon. The calculations are performed with the nucleon carrying different momenta $P_{z}^{n}=(2\pi n)/L$ along the $z$ direction, and the total momentum of the dibaryon system is $2P_{z}^{n}$ For each momentum, several sets of source-sink separations $t_{sep}$ are calculated. For larger momenta, correlators are averaged over several sources evenly distributed on the $z=0$ plane to control the statistical uncertainties. The detailed information is shown in Table 2.
Ensemble $P_{z}^{n}$(GeV) $t_{sep}/a$ $N_{conf}$ $nsrc(time\times space)$
@@
Table 2: Information on the nucleon’s momentum $P_{z}^{n}$ (the center-of-mass momentum of the dibaryon system is $2P_{z}^{n}$), time separation between source and sink($t_{sep}$) and number of sources($nsrc$) on the $z=0$ plane used to calculate two-point and three-point correlators.
To improve the signal of high-momentum states, we applied momentum-Gaussian smearing on sources and sinks Bali et al. (2016), which takes the form
@@
where $k$ represents the momentum smearing parameter and $U_{j}$ denotes a stout-smeared gauge link along a specified direction. The above operation should be iterated by $n_{iter}=\mathcal{O}(2\sigma^{2})$ times to achieve the desired smearing size $\sigma a$, which corresponds to $n_{iter}=50$ and $\sigma a=0.5$ fm in our calculation. This process shifts the smearing center to momentum $\mathcal{O}(k)$, thereby significantly improving the signal of high-momentum states. In our calculation we set $k$ equal to half of the nucleon momentum, which is the optimized choice to improve the signal according to previous experience Bali et al. (2016). In addition, we applied one-step hypercubic(HYP) smearing on the Wilson link of the current operator.